# T-MAE: Temporal Masked Autoencoders for Point Cloud Representation Learning

### v1

@@ 3.2 Temporal Masked Autoencoder Pre-training
Inspired by the SiamMAE [17] used in video understanding, we develop Temporal Masked Auto-Encoders, namely T-MAE, for self-supervised learning on LiDAR points. As shown in Fig. 2, the core idea is to reconstruct the present frame based on the historical and partial observations of the current frame. In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion. After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks. In the subsequent subsections, we provide a detailed elaboration of these steps.
Single-frame baseline.
To perform the reconstruction pretext task, we follow the concept of the state-of-the-art SSL method GD-MAE [52] which operates on a single frame. The pipeline encompasses several key stages, including voxelization, masking, encoding, decoding, and reconstruction. More specifically, the initial step voxelizes a given point cloud by two linear layers and a voxel-wise average pooling. The voxel representation can be considered a pillar-based representation, owing to its infinite height. Next, masking is applied in a pillar-wise manner. Subsequently, the remaining visible pillars serve as input for the sparse pyramid transformer to generate high-level tokens. Next, the sparse tokens are transformed into a dense tensor by means of zero-padding. Subsequently, a series of dense convolutional layers are utilized to dilate the tokens of unmasked voxels to vacant regions.
$t$ ${t-3}\ \sim\ t$
Refer to caption Refer to caption
(a) A stationary vehicle.
Refer to caption Refer to caption (b) A moving vehicle.
Figure 3: Comparison between single and four frame concatenation. While simple frame concatenation generally improves point density and detection rates, it can introduce spurious points in non-static scene parts that may degrade the detection performance. Since we combine consecutive frames via learned cross-attention, our approach is less affected by this problem. The blue bounding boxes indicate the ground truth for the current frame.
@@
Figure 4: Window-based sparse cross-attention (WCA) module. Given the input tokens from both the previous and current frames, a joint region grouping is conducted to obtain a window partition. A sparse regional cross-attention (SRCA) is performed individually in each window to integrate the historical information to the middle tokens for the current frame. In other words, the tokens from two frames but with the same colors are attending to each other. For simplicity, the information flow is only illustrated for the green tokens. The grouping and cross-attention are performed once more with the shifted window partition. The red dot (Refer to caption) indicates the ego-vehicle driving towards the right. The box with diagonal stripes (Refer to caption) represents an object moving towards the left. The observed scene remains mostly static except for a pillar at bottom-left transitions from occupied to vacant due to occlusion and another pillar which does the opposite. Best viewed in color and high-resolution.
SiamWCA.
Up to this point, a common approach has been to employ a single-frame reconstruction pretext task for LiDAR points. However, incorporating historical frames poses a non-trivial challenge. One straightforward approach is to concatenate two point clouds after aligning them with ego-poses, similar to the three-frame variant of SST [13]. However, as shown in Fig. 3, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects. We provide quantitative results in Tab. 2. To address this problem, we suggest learning from past data in a feature space.
Given two frames, the previous frame is transformed to the coordinate system of the later frame by their ego-poses. Pose information is usually readily available in relevant datasets [39, 2, 29] or easily obtained by means of GPS/IMU, odometry approaches [8], structure from motion (SfM) algorithms [37, 38], or SLAM systems [10]. As shown in Fig. 2, the full observation of the previous frame is given to the network by passing it through voxelization and encoder. On the contrary, the later frame goes through voxelization, masking, and encoder. The encoders for the two frames share weights to ensure that the features are constrained in the same space.
By utilizing the Siamese encoder, tokens are obtained for both frames. These tokens serve as input to the WCA module, which facilitates the interaction between historical and current tokens and is outlined in Sec. 3.3. Once the current tokens are augmented with historical information, these sparse tokens are placed back into the x-y plane to form a dense tensor where vacant pillars are padded with zeros. Subsequently, a group of $3\times 3$ convolutional layers is utilized to process the tensor, which diffuses the features of visible pillars. Eventually, the obtained tokens of masked pillars are employed for point reconstruction. The Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points. Note that the number of points per pillar varies drastically. Thus, a fixed number of points $K^{GT}$ are randomly sampled as the target for reconstruction.
In conclusion, SiamWCA, containing a Siamese encoder and a WCA module, computes the pillar-based representation for both frames and integrates the historical one to the present one for reconstruction. Unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks.
3.3 Window-based Sparse Cross-Attention Module
When provided with input tokens from two frames, there are several methods for interaction. One intuitive approach is to apply a standard self-attention layer to the concatenation of all tokens from both frames. This approach significantly increases GPU memory requirements. The vanilla Transformer block [43], which is combined with a cross-attention layer and a self-attention layer, can also be adopted. This combination, however, is demonstrated to be suboptimal as outlined in Sec. 4.3. In due course, a single cross-attention layer is evaluated and exhibits superior performance compared to others.
However, conventional global attention is not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution. Therefore, the use of window-based attention is vital. The WCA module is illustrated in Fig. 4. The simulation involves an ego-vehicle moving to the right and another object moving in the opposite direction. The scene remains mostly static, but a pillar at bottom-left transitions from occupied to vacant due to occlusion.
Joint Regional Grouping
organizes the 3D spaces into non-overlapping windows. Hence, the following attention mechanism only performs with tokens from the same local region. The tokens from two frames are jointly considered so that the local windows across the two frames are corresponding. As shown in Fig. 4, all tokens within the same physical window share the same color, indicating that they are assigned to the same group for mutual attention.
Sparse Regional Cross-Attention
@@ 3.3 Window-based Sparse Cross-Attention Module
(SRCA) is essentially a cross-attention layer where the query comes from the current frame and the key-value comes from the previous frame. For a group of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ are calculated as follows
$\mathcal{Q}$ $=\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}})$ (1)
$\mathcal{K}$ $=\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}})$ (2)
$\mathcal{V}$ $=\mathcal{F}^{t_{1}}$ (3)
where $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [4]. SRCA follows conventional cross-attention and is defined as follows
^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ (4)
~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)$ (5)
where $\mathbf{MCA}(\cdot,\cdot,\cdot)$ indicates the Multi-head Cross-Attention, and $\mathbf{LN}(\cdot)$ stands for Layer Normalization. While the core concept is similar to cross-attention, this implementation provides a more efficient solution when working with sparse pillars.
Repeated Operations with Region Shift.
The joint regional grouping and SRCA are performed once more after the regions are shifted. We follow the implementation of SST [13] to achieve region shift. Formally, suppose that the window size is $(l_{x},l_{y},l_{z})$, then the original windows are moved by $(l_{x}/2,l_{y}/2,l_{z}/2)$. Meanwhile, the tokens are re-grouped by the joint regional grouping, as illustrated in “Shifted Window Partition” of Fig. 4. Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of the current frame are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged.
@@ 4.1 Dataset and Implementation
Waymo Open dataset [39] is the currently the largest public dataset with LiDAR points for autonomous driving. It comprises 798 sequences for training and 202 sequences for validation. Each sequence contains roughly 200 frames with pose information by which points can be transformed into the world coordinate system. For 3D detection, an evaluation protocol is provided for calculating the average precision (AP) and the average precision weighted by heading (APH). Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1 while Level 2 indicates all bounding boxes. We adopt mean AP and mean APH at Level 2 as main evaluation metrics.
Implementation Details. We implement our approach based on the codebase of OpenPCDet [^1] https://github.com/open-mmlab/OpenPCDet. We keep the configuration of the sparse pyramid transformer consistent with its official implementation of GD-MAE [52]. The window size of the WCA modules is set to $(8,8,1)$. The detection ranges for the Waymo dataset are restricted between $(-74.88m,-74.88m,-2m)$ and $(74.88m,74.88m,4m)$. The pillar size is $(0.32m,0.32m,6m)$. The size of the temporal batch is set to $n=6$ for both pre-training and finetuning. Three data augmentation techniques, i.e. random flipping, scaling, and rotation, are applied to both frames during pre-training and finetuning. Note that, the variables of augmentations applied to two frames keep the identical. We use the AdamW optimizer [28] with the one-cycle policy to train the model. The max learning rate is set to $3e^{-3}$. All experiments are performed using 8 NVIDIA Tesla V100.
Pre-training. The model is pre-trained for 12 epochs on the full training split of Waymo dataset, which contains 158,081 frames. The batch size is set to 4. Following GD-MAE [52], the masking ratio is set to 0.75, the number of predicted points per pillar is set as 16 and the number of target reconstruction points is set as $K^{GT}=64$.
Finetuning and inference. The model is finetuned for 30 epochs on different subsets of the Waymo training split. The number of epochs is held constant at 30, with a batch size of 2, while the data quantity varies. In addition to the data augmentations mentioned in the implementation details of Sec 4.1, following MV-JAR [48] and GD-MAE [52], a copy-n-paste augmentation [51] is also employed to slightly address the issue of class imbalance. During inference, each frame is selected as a current frame for a complete evaluation. The previous third frame, based on our ablation study, is selected as the previous frame.
4.2 Results on Waymo
In this section, the main results of the Waymo dataset are presented. We assess our approach after fine-tuning SiamWCA with varying amounts of labeled data. We explore two methods to obtain a specific percentage of the training set, but all models are evaluated on the same validation set.
$\triangleright$ Uniform sampling is commonly used to obtain a subset of LiDAR sequence [53, 52, 25]. Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval.
@@ 4.2 Results on Waymo
L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
@@
GD-MAE [$*{}^{*}$ 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
T-MAE (Ours) 63.48 ${}^{+3.27}$ 61.72 ${}^{+4.11}$ 62.93 62.47 68.26 64.53 59.24 58.16
ProposalContrast [53] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
BEV-MAE [26] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
Occupancy-MAE [31] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
MAELi-MAE [22] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
GD-MAE [${}^{**}$ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
AD-PT [55] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
T-MAE (Ours) 70.61 ${}^{+1.58}$ 68.80 ${}^{+2.32}$ 66.85 66.39 74.85 71.08 70.14 68.94
@@
GD-MAE [${}^{**}$ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
T-MAE (Ours) 71.96 ${}^{+0.65}$ 70.22 ${}^{+1.09}$ 69.07 68.63 74.74 71.03 72.06 70.99
Table 1: Performance comparison with SSL methods on the Waymo validation set [39]. Random initialization denotes training frame scratch. 20% Uniform Sampling indicates sampling frames per five frames. $*{}^{*}$ indicates reproduced by us. ${}^{**}$ indicates from [52]. Results for other methods are taken from MV-JAR [48] or the survey [14]. Best results are highlighted as first, second, and third. Differences between T-MAE pre-training and random initialization are highlighted in red. T-MAE achieves significant improvements compared to random initialization and SOTA results compared to other SSL methods.
$\triangleright$ Data-efficient benchmark is proposed in MV-JAR [48] to solve the data diversity issue existing in the strategy of uniform sampling. More specifically, while trained on uniformly sampled frames, the models always converge to a similar performance if they are trained for adequate iterations no matter how many percentages of labeled data are used. As a result, fewer data do not lead to a shorter finetuning time. Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab. 1. Note that, for each data amount level, three non-overlap splits are provided MV-JAR [48]. The results for models finetuned by the second or the third split are presented in the supplementary material, as well as the average performance.
$\triangleright$ T-MAE achieves new SOTA at labelled data level of 10%, 20%, 20% (Uniform sampling) and 100%. Moreover, T-MAE achieves superior performance for the majority of class-specific metrics. To reduce performance variance on 5% and 10% splits, we also report in the supplementary material the results of T-MAE by finetuning it on the other two splits provided by [52]. Notably, thanks to the temporal modeling ability, T-MAE significantly outperforms other methods in terms of L2 mAPH for pedestrians. For instance, with 5% labeled data, T-MAE achieve better mAPH (i.e. 55.71) than any other method using 10% labeled data. The comparison with other SSL methods shows the efficacy of T-MAE. Nevertheless, MV-JAR [48] performs notably better for cyclists at 5% level and surpasses T-MAE in terms of L2 overall mAP.
@@
$\triangleright$ The impact of the T-MAE pre-training becomes more pronounced when the labeled data is scarce. SiamWCA is a powerful backbone as, when finetuned with 100% of labeled data, it even performs better than other models that are initialized by different pre-training strategy (i.e. 69.13 v.s. 67.64). However, its performance drops significantly when finetuning data is limited (i.e. 40.29 v.s. 46.68 at 5% level). T-MAE consistently enhances model performance compared to random initialization, with more significant improvements observed when labeled data is limited.

### v3

@@ 3.3 Temporal Masked Autoencoder Pre-training
Up to this point, a two-frame framework for object detection has been set up. However, the transformer-based architecture is data-hungry. Inspired by the SiamMAE [22] used in video understanding, we develop Temporal Masked Auto-Encoders (T-MAE) for self-supervised learning on LiDAR points. As shown in Fig. 3, the core idea is to reconstruct the present frame based on a full observation of the historical frame and a partial observation of the current frame. In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion. After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks. In the subsequent paragraphs, we elaborate on these steps.
Masking. As shown in Fig. 3, on top of the proposed SiamWCA backbone, an additional step, namely masking, is added between the voxelization and the encoder for the current frame. Specifically, masking is applied in a pillar-wise manner for a high ratio of the occupied pillars, e.g., 75%. The remaining pillars are subsequently input into the encoder. The encoders for the two frames share weights to ensure that the features are constrained in the same latent space. Due to masking, the number of tokens for the current frame decreases significantly, leading to much fewer valid windows and thus accelerating WCA during pre-training. The dense feature recovery performs exactly the same as in Sec. 3.2 and outputs a dense feature map.
Reconstruction head. Given the dense feature map, the reconstruction head retrieves the feature of masked pillars by their spatial location. With the pillar-wise features, the head reconstructs the relative coordinates of points, where the number of output points per pillar is set as $K^{O}$. Eventually, the Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points. Note that the number of points per pillar varies drastically. Thus, a fixed number of points $K^{GT}$ are randomly sampled as the target for reconstruction. In conclusion, T-MAE masks the voxelized tokens of the current frame and compels the network to reconstruct the current frame with the full observation of the historic frame as a reference, which encourages the WCA module to build up correspondence between frames. Note that, unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks, which is proved to be effective as shown in Tab. S10. Therefore, the capability for building up correspondence is also retained.
3.4 Siamese Encoder and Windowed Cross-Attention (SiamWCA)
@@
Figure 4: Windowed sparse cross-attention (WCA). Given the input tokens from both $\mathcal{P}^{t_{1}}$ and $\mathcal{P}^{t_{2}}$, a joint token grouping is performed to obtain a window partition. A sparse regional cross-attention (SRCA) is performed independently in each window to integrate the historical information to the middle tokens of the current frame. In other words, the tokens from two frames but with the same colors are attending to each other. For simplicity, the information flow is only depicted for the green tokens. After the second joint token grouping, the cross-attention are performed once more with the shifted window partition. The red dot (Refer to caption) indicates the ego-vehicle driving towards the right. The box with diagonal stripes (Refer to caption) represents an object, e.g., a vehicle, moving towards the left. Best viewed in color and high-resolution.
Siamese encoder. Given the pillar-wise representations of a pair of frames, we explore an asymmetric network and a Siamese encoder for feature encoding. The asymmetric network consists of two branches with the same architecture but the branch for $\mathcal{P}^{t_{1}}$ is modified by reducing the number of channels by half, as depicted in Fig. 6 (a). A Siamese encoder is a symmetric network with two subnetworks sharing weights. It is widely used to compute similarity in latent space for tracking [1, 23, 51, 21] and contrastive learning [10, 44, 12]. For the Siamese encoder, we investigate two weight updating strategies, namely accumulation and SimSiam-style [10]. Accumulation indicates the backward gradients of the two encoders are accumulated. SimSiam, standing for the simple Siamese pre-training strategy [10], indicates the encoder for $\mathcal{P}^{t_{1}}$ is detached from the computational graph, as depicted in Fig. 6 (b). Thus the weight of this encoder is not updated by gradient propagation but by copying the weights from the encoder for $\mathcal{P}^{t_{2}}$.
Windowed cross-attention (WCA). When provided with input tokens from two frames, there are several methods for interaction. One intuitive approach is to apply a standard self-attention layer to the concatenation of all tokens from both frames. This approach significantly increases GPU memory requirements because it doubles the number of input tokens. The vanilla Transformer block [55], which consists of a cross-attention layer and a self-attention layer can also be adopted and a single cross-attention layer is feasible as well. However, all these conventional global attentions are not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution. For instance, the input for our Siamese encoder can be $\mathbb{R}^{(468\times 468)\times 128}$, whereas it is $\mathbb{R}^{(14\times 14)\times 384}$ for a ViT-B/16 due to the patch embedding. Therefore, a window-based implementation is crucial for efficiency. In brief, the WCA module divides the 3D space into non-overlapping windows and then performs cross-attention within the windows. Figure 4 intuitively illustrates this process with a diagram that simulates a scene from a bird’s eye view. The key components are elaborated in the following paragraphs.
$\triangleright$ Joint token grouping partitions the 3D space into non-overlapping windows and subsequently allocates the token to the corresponding window based on its spatial location. The previous frame has been aligned with the current frame and thus the window partition is unified for both frames, which allows the following attention mechanism to perform within each window. As shown in Fig. 4, all tokens within the same physical window share the same colour, indicating that they are assigned to the same group for mutual attention.
@@
$\triangleright$ Sparse Regional Cross-Attention (SRCA) is essentially a cross-attention layer where the query comes from $\mathcal{P}^{t_{2}}$ and the key-value comes from $\mathcal{P}^{t_{1}}$. For clarity, we consider a single window as an example. Given two groups of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ from two frames and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the cross-attention is performed as follows:
^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}\left(\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}}),\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}}),\mathcal{F}^{t_{1}}\right)$ (1)
~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)+\mathcal{F}^{t_{2}}$ (2)
where $\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ indicates a classical Multi-head Cross-Attention, $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [6], and $\mathbf{LN}(\cdot)$ stands for Layer Normalization. Note that, if a window is empty in the previous frame, the tokens of this window in the current frame will remain unchanged, i.e. ~$\widetilde{\mathcal{F}}^{t_{2}}=\mathcal{F}^{t_{2}}$. While the core concept of WCA is identical to cross-attention, this windowed implementation significantly scales down the computational complexity especially when working with sparse pillars.
$\triangleright$ Repeated operations with window shift. Given the middle tokens, the window partition is shifted by half of the window size. Next, the tokens are re-grouped by the joint token grouping, as illustrated in “Shifted Window Partition” of Fig. 4. After that, the SRCA is performed once more. Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of $\mathcal{P}^{t_{2}}$ are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged because a cross-attention operation does not update key $\mathcal{K}$ and value $\mathcal{V}$. With this window shift and SRCA #2, a token interacts with more tokens. For instance, the yellow token in the first window partition only attaches itself in the previous frame via SRCA #1 but interacts with more tokens in SRCA #2.
@@ 4.1 Dataset and Implementation
Experiments are conducted on the following two datasets.
Waymo Open dataset [49] is a large-scale autonomous driving dataset with LiDAR points. For 3D detection, an evaluation protocol is provided for calculating the average precision (AP) and the average precision weighted by heading (APH). Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1, while Level 2 indicates all bounding boxes. We adopt mean AP and mean APH at Level 2 as the main evaluation metrics.
ONCE dataset [37] consists of 581 sequences of varying length. 6, 4 and 10 sequences are selected for training, validation and test sets, respectively, and manually annotated. The remaining data is not annotated and adopted for pre-training in our experiments. The official evaluation metrics are AP calculated for each category and range.
Implementation Details. We implement our approach based on the codebase of OpenPCDet [^1] https://github.com/open-mmlab/OpenPCDet. We adopt the sparse pyramid transformer as our encoder and keep the configuration consistent with its official implementation of GD-MAE [64]. Three data augmentation techniques, i.e. random flipping, scaling, and rotation, are applied to both frames during pre-training and finetuning. In addition, following MV-JAR [60] and GD-MAE [64], a copy-n-paste augmentation [63] is also employed to slightly address the issue of class imbalance during finetuning. Note that identical data augmentations are applied to a pair of two frames. Further details are in the supplementary material.
4.2 Main Results
We present the comparison with SOTA methods on two datasets: For the ONCE dataset [37], we pre-train our SiamWCA with the raw_large split and fine-tune it with the annotated training set. For the Waymo dataset [49], we pre-train SiamWCA with the entire training set. Then, following MV-JAR [60], we finetune our approach with four portions of labeled data, namely 5%, 10%, 20% and 100%. Note that, a strong counterpart, GD-MAE [64], has not been evaluated in this setting and thus we re-train it for a more comprehensive comparison.
[…]
@@ 4.2 Main Results
mAP mAPH mAP mAPH mAP mAPH mAP mAPH
@@
GD-MAE [64] 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
T-MAE 62.37 60.17 62.19 61.72 67.18 62.18 57.74 56.59
T-MAE (Ours) 63.52↑3.31 61.80↑4.19 63.10 62.59 68.23 64.66 59.23 58.15
@@
GD-MAE [64]∗∗ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
T-MAE 71.56 69.00 69.39 68.95 74.42 68.43 70.86 69.61
T-MAE (Ours) 72.30↑1.00 70.52↑1.39 69.34 68.89 75.79 72.01 71.78 70.65
@@
The impact of the T-MAE pre-training. The bottom block in Tab. 1 shows that the randomly initialized SiamWCA (denoted as Random) performs better than any other models that are initialized with a different pre-training strategy (i.e. 69.13 v.s. 67.64), suggesting that SiamWCA is a powerful backbone capable of learning temporal modeling when provided with sufficient annotated data. However, its performance drops significantly when finetuning data is limited (e.g., 40.29 v.s. 46.68 at 5% level), indicating a strong demand for annotated data. As a comparison, the proposed T-MAE consistently enhances SiamWCA compared to random initialization. Moreover, as the labeled data shrinks from 100% to 5%, the impact of the T-MAE pre-training becomes more pronounced, namely increasing from 1.39 to 9.17, suggesting that the pre-training approach learns a powerful representation and alleviates the demand for annotated data.
Figure 5: Qualitative results. We depict ground truth and predictions as boxes colored in red and green for two exemplary scenes from the Waymo dataset [49].
Comparison with SOTA methods. We aim to leverage the temporal information between two adjacent frames, which is often overlooked by other methods. This absence makes it challenging to compare our method with others in the same setting. Therefore, we implement a test-time single-frame baseline (denoted as T-MAE) by replicating the same frame and inputting them into our pre-trained model during evaluation. Table 1 shows that T-MAE with identical frames outperforms SOTA counterparts. Moreover, T-MAE with adjacent frames achieves new SOTA at all levels in terms of overall and class-specific metrics. Notably, thanks to the temporal modeling ability, T-MAE significantly outperforms other methods in terms of L2 mAPH for pedestrians. For instance, with 5% labeled data, T-MAE achieve better mAPH (i.e. 55.28) than any other method using 10% labeled data. This metric indicates better direction detection for pedestrians, which could benefit downstream applications, e.g., pedestrian intention prediction.
To reduce performance variance on 5% and 10% splits, we also report T-MAE performance in the Appendix (see Sec. S4) when it is finetuned with two more 5% and 10% splits and another 20% split. T-MAE consistently outperforms other methods, indicating the efficacy of T-MAE. Since there are no existing multi-frame SSL methods, we also compare our approach with robust non-SSL baselines that utilize multi-frame as inputs in the Appendix (see Tab. S9).
To assess the generalization capabilities of our method, we also conducted experiments on the ONCE dataset [37]. As shown in Tab. 2, T-MAE outperforms other methods in most metrics, indicating its superiority. Moreover, the substantial improvement for pedestrians generalizes to this new dataset, indicating the dominance of T-MAE in pedestrian detection.
Figure 5 shows two exemplary qualitative results from the Waymo dataset and more qualitative results are presented in the Appendix.
Table 2: Performance comparisons on the validation split of the ONCE dataset [37]. Pt. indicates the model is initialized with pre-trained weights. Results for other methods are taken from GD-MAE [64].
Methods Pt. mAP Vehicle Pedestrian Cyclist
Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf
PV-RCNN [48] 53.55 77.77 89.39 72.55 58.64 23.50 25.61 22.84 17.27 59.37 71.66 52.58 36.17
IA-SSD [73] 57.43 70.30 83.01 62.84 47.01 39.82 47.45 32.75 18.99 62.17 73.78 56.31 39.53
CenterPoint-Pillar [68] 59.07 74.10 85.23 69.22 53.14 40.94 48.43 34.72 20.09 62.17 73.70 56.05 40.19
CenterPoint-Voxel [68] 60.05 66.79 80.10 59.55 43.39 49.90 56.24 42.61 26.27 63.45 74.28 57.94 41.48
SECOND [63] 51.89 71.19 84.04 63.02 47.25 26.44 29.33 24.05 18.05 58.04 69.96 52.43 34.61
w/ BYOL [20] 51.63 71.32 83.59 64.89 50.27 25.02 27.06 22.96 17.04 58.56 70.18 52.74 36.32
w/ PointContrast [59] 53.59↑1.70 71.87 86.93 62.85 48.65 28.03 33.07 25.91 14.44 60.88 71.12 55.77 36.78
w/ DeepCluster [52] 53.72↑1.83 72.89 83.52 67.09 50.38 30.32 34.76 26.43 18.33 57.94 69.18 52.42 34.36
SPT [64] 62.62 75.64 87.21 70.10 53.21 45.92 54.78 37.84 22.56 66.30 78.12 60.52 42.05
w/ GD-MAE [64] 64.92↑2.30 76.79 88.01 71.70 55.60 48.84 58.70 37.30 25.72 69.14 80.29 64.58 45.14
SiamWCA (Ours) 63.71 76.47 87.63 71.59 55.16 47.27 57.57 36.99 21.79 67.40 78.39 62.78 43.90
w/ T-MAE (Ours) 67.00↑3.29 78.35 88.45 73.05 57.16 52.57 62.66 44.18 25.29 70.09 81.14 65.33 46.48